Find the present value of $10,000 received at the start of every year for 20 years if the interest rate is J1 = 12% p.a. and if the first payment of $10,000 is received at the end of 10 years
1
Expert's answer
2020-06-14T17:27:34-0400

The present value will be calculated as follows:

The interest for the 10 years:

"\text{Simple interest} = 10000\times10\times \dfrac{12}{100}= 12000"

Therefore, the amount will be:

"\text{Amount}=10000+12000=22000"

after 20yrs another payment will be received value of principal is 10000*2 = 20000

The interest will therefore, be:

"\\text{Simple interest}=20000\u00d720\u00d7\\dfrac{12}{100}\u200b=48,000"


"\text{Amount}=48,000+20,000=68,000"

Present value will therefore be:

"\text{Present value}=22000\times \dfrac{1}{(1+0.12)^{10}}+68000\times \dfrac{1}{(1+0.12)^{20}}"



"\text{Present value}=14,132.75"